OpenStudy (anonymous):

Using the equation, y = 12841(1.01)^x, predict the cost of an education in a private college in the year 2010. Use x = 20 as this is 20 years after 1990. Round to the nearest year.

OpenStudy (anonymous):

plug in 20 for x: y=12841(1.01)^20 y=12,841*1.22019 y=15,668